Alice is 10 years older as her brother Alex. In 5 years, she will be twice as old as her brother. Find the present ages of each

Answer by Cromlix(4381)

Alice is 10 years older as her brother Alex.
Al = 10 + Ax
In 5 years, she will be twice as old as her brother.
Al + 5 = 2(Ax + 5)
Replace Al with 10 + Ax from top equation.
10 + Ax + 5 = 2Ax + 10
Ax - 2Ax = 10 - 10 - 5
-Ax = -5
Ax = 5
Alex is 5 years old
Alice is 15 years old.
Hope this helps.
:-)